# Error Crafting in Probabilistic Quantum Gate Synthesis

### v1

@@ Introduction
If one choose to employ the PEC method as the QEM method for FTQC, it suffices to guarantee only the feasibility under the Pauli constraint. Indeed, the PEC method is powerful in the sense that one may perform unbiased estimation when the error channel is precisely known. However, we find that the remnant errors can be counteracted much more effectively by utilizing the fact that they can be characterized completely by classical computation. For instance, we may wish to craft local depolarizing errors, since many limits and performance of quantum algorithms as well as QEM methods are analyzed under the assumption of the local depolarizing noise [32, 19, 52, 35, 53]. We may alternatively wish to craft the remnant errors to be some biased Pauli noise. We find that this is extremely beneficial for case of the Pauli rotation gates. Namely, we can perform error detection or correction in the logical layer, which reduces the sampling overhead to the extent that QEM methods designed for general channels cannot achieve.
@@
First, let us discuss the crafting of the depolarizing channel. Following a similar strategy as in Theorem 1, we can mathematically prove that there exists a set of shift unitaries $\{\mathcal{V}^{c\epsilon}_{j}\}$ such that we can achieve quadratic suppression of error under the depolarizing constraint, with the sole difference being that the number of shift unitary basis is 9 instead of 7. The formal statement and the proof are provided in Supplementary Note S6. Our numerical simulation also shows similar favorable properties as seen in the case for Pauli constraint (See Supplementary Note S2).
@@
Next, let us consider crafting the remnant error to be biased. In particular, we focus on the Pauli Z rotation gate and aim to bias the error to be XY error, since it can be detected using a single ancilla qubit with less sampling overhead than the lower bound for QEM methods that does not rely on mid-circuit measurement (concrete circuit structure provided in Fig. 3). While we have not found a guaranteed set of shift unitary as in the case for the depolarizing noise, we employ a general strategy of expanding the size of basis set until the synthesis is successful, i.e., the contribution of Z error is below a fixed threshold value (see Supplementary Note S2)
@@
The success of error crafting is highlighted in Fig. 2. As can be seen from Fig. 2(a), while the rate of Pauli errors $(\frac{p_{x}}{p},\frac{p_{y}}{p},\frac{p_{z}}{p})$ is distributed quite homogeneously if we do not impose any constraint on the error profile, the constraints can be reflected with high accuracy. We verify that $p_{x}=p_{y}=p_{z}$ is satisfied up to order of machine precision for the depolarizing constraint, while for the XY-biased noise, we find that we cannot completely eliminate the Z error; there is a small residual component of $p_{z}=O(\epsilon^{3}).$
@@
Figure 3: Circuit structure for (a) error detection and (b) logical error correction for Pauli Z rotation gate $R_{z}$ with remnant error of $\mathcal{E}_{\rm rem}$. We may further combine with the PEC method, for which the Pauli operation $\mathcal{P}$ with dotted square is executed virtually via updating the Pauli frame. Note that the feedback operation in (b) is also done via Pauli frame.
@@
Gate Synthesis Remnant error QEM method Sampling overhead Novelty Comment
@@
SU(2) Crafted $\epsilon^{2}$ Pauli/depol PEC $1+4\epsilon^{2}$
SU(2) Crafted $\epsilon^{2}$ Pauli/depol Rescaling $1+2\epsilon^{2}$
$R_{z}$ Unitary $\epsilon$ coherent PEC $1+4\epsilon$
@@
$R_{z}$ Crafted $\epsilon^{2}$ Pauli/depol/XY PEC $1+4\epsilon^{2}$
$R_{z}$ Crafted $\epsilon^{2}$ Pauli/depol/XY ED + PEC $1+(q_{x}+q_{y}+4q_{z})\epsilon^{2}$ $q_{z}\sim\epsilon$ in numerics
$R_{z}$ Crafted $\epsilon^{2}$ Pauli/depol/XY ED + Resc. $1+(q_{x}+q_{y}+2q_{z})\epsilon^{2}$ $q_{z}\sim\epsilon$ in numerics
$R_{z}$ Crafted* $\epsilon^{3}$ Pauli Z PEC $1+4\epsilon^{3}$
$R_{z}$ Crafted* $\epsilon^{3}$ Pauli Z Rescaling $1+2\epsilon^{3}$
@@
Table 1: Sampling overhead of error countermeasures under synthesis error of $\epsilon$. The remnant error indicates the diamond distance $d(\mathcal{E}_{\rm rem},\mathcal{I})$ where $\mathcal{I}$ is the identity channel. We assume that unitary gate synthesis is done with accuracy of $\epsilon$, and compare the performance when we further perform probabilistic or error-crafted synthesis. The sampling overhead $\gamma$ is indicated for a single gate, and the total overhead is multiplicative as $\prod_{l}\gamma_{l}$ when overhead is $\gamma_{l}$ for the $l$-th gate. The rescaling technique assumes that the underlying quantum circuit is sufficiently deep so that the white noise (WN) approximation holds. We assume that the white noise approximation holds when rescaling technique is applied. The cubic suppression in the last two rows are verified numerically up to $\epsilon^{3}=10^{-9}$ as shown in Fig. 4. Note that * for the last two rows indicate that synthesis is accompanied with a single ancilla qubit.
@@
Figure 4: Trade-off between T-count and synthesis accuracy for Pauli rotation gates. We assume that each T gate is subject to depolarizing error of rate $p=10^{-10}$ while Clifford gates are noiseless. The red, yellow, and blue regions indicate the achievable accuracy without any sampling overhead when we employ unitary, probabilistic, and error-crafted synthesis. The blue circuits are the results obtained for $R_{z}(\theta)$ with $\theta\in\{\frac{m\pi}{32}\}_{m\neq 0~{}{\rm mod4}}$ under various synthesis accuracy. The dotted red, yellow, and green lines are given by the empirical T count of $3\log_{2}(1/\varepsilon)$, $1.5\log_{2}(1/\varepsilon)$, and $\log_{2}(1/\varepsilon)$. The chained black line is given by the accumulation of incoherent noise in the T gates.
@@
Even if the Z error cannot be eliminated completely, such an error crafting drastically alters the sampling overhead. By performing the error detection operation as shown in Fig. 3(a), we can mitigate Pauli X and Y errors with sampling overhead of $\gamma_{\rm XY}=1+p_{x}+p_{y},$ which is even below the lower bound of QEM sampling overhead that does not rely on mid-circuit measurement [33]. While the Z error is undetectable and hence we shall rely on other QEM methods such as PEC or rescaling with overhead of $\gamma_{Z}=1+\alpha p_{z}$, the total sampling overhead is $\gamma_{\rm tot}=\gamma_{\rm XY}\gamma_{Z}\sim 1+p_{x}+p_{y}+\alpha p_{z}\sim 1+\epsilon^{2}$. When we perform QEM for $L$ gates, the total sampling overhead $\gamma_{\rm tot}^{L}$ is quartically small compared to that of PEC.
Error crafting using CPTP maps
While we have considered the error crafting using synthesized unitaries $\{U_{j}\}$, we can also apply the framework to CPTP maps. As a demonstration, here we consider the target unitary $\hat{U}$ to be a Pauli Z rotation, and constitute the set of implementable CPTPs $\Lambda_{j}$ as the synthesized unitary accompanied with entangled measurement and feedback as shown in Fig. 3(b). By performing brute-force search of $\{\Lambda_{j}\},$ we find that the remnant error can be suppressed cubically small as $\epsilon^{3}$, as opposed to the quadratic suppression realized in the error crafting for unitaries. As we show in Fig. 4, this allows us to synthesize a Pauli rotation gate with T count of $\log_{2}(1/\varepsilon)$ on average, up to accuracy of $\varepsilon=10^{-9}.$ We guide the readers to the Supplementary Note S3S4 for detailed discussion on the search for appropriate $\Lambda_{j}$.
@@
Let us briefly comment on the classical precomputation time of crafted synthesis. Our results in Fig. 4 exploits the direct search algorithm for SU(2) gate synthesis whose computational complexity scales as $O((1/\epsilon))$ on average [51]. Considering that currently known unitary synthesis algorithms for $R_{z}$ gates based on number-theoretic approach [54, 42] or repeat-until-success circuits [46, 39] yield nearly optimal gate sequence with T count of $O({\rm polylog}(1/\epsilon))$, the scaling of direct search algorithms seems to be a significant overhead. Meanwhile, we observe that the actual computation time per synthesis is observed to be milliseconds for the accuracy regime we have investigated. Considering that such computations can be embarrassingly parallelized, and also that number of rotation gates for early FTQC circuits is of $10^{4}$ for quantum dynamics simulation [55, 30] and $10^{6}$ for ground state energy simulation [56], we envision that classical computation time is not bottleneck for such applications.
@@
Last but not least, while we have exclusively discussed the error and sampling overhead only for the synthesis error, it must be kept in mind that, in reality, there is also incoherent noise that indirectly limits the synthesis accuracy. As can be seen from Fig. 4, one shall aim for the sweet-spot where the total error is minimized. While we have displayed the case when we have designed the FTQC to contribute the majority of logical error to magic state distillation, we may flexibly change the scheme to introduce errors in Clifford operations as well.
Summary and outlook
@@
Numerous future directions are envisioned. First, it is essential to seek whether the error crafting is possible for synthesis of multiqubit unitaries. Combined with the fact that the error can potentially be suppressed exponentially smaller with the qubit count for multiqubit case [41], it is crucial to extend the framework proposed in this work to perform practical quantum computation. Second, it is intriguing in terms of resource theoretic argument to assess the probabilistic transformability between channels. While this has been studied for state transformation in the context of magic state or entanglement distillation [57], the argument on quantum channels is left as an interesting open question. Third, it is an intriguing open question whether there is lower bound on the computational complexity to perform error-crafted synthesis.
@@ Acknowledgement
We thank fruitful discussions with Suguru Endo, Yuuki Tokunaga, and Kunal Sharma. This work was supported by JST CREST Grant Number JPMJCR23I4, JST Moonshot R&D Grant Number JPMJMS2061, MEXT Q-LEAP Grant No. JPMXS0120319794, and No. JPMXS0118068682. N.Y. wishes to thank JST PRESTO No. JPMJPR2119, JST ERATO Grant Number JPMJER2302, JST Grant Number JPMJPF2221, and the support from IBM Quantum. S.A. is supported by JST, PRESTO Grant no.JPMJPR2111. This work is supported by MEXT Quantum Leap Flagship Program (MEXT Q-LEAP) Grant No. JP- MXS0118067394 and JPMXS0120319794, JST COI- NEXT Grant No. JPMJPF2014. K.T. is supported by World-leading Innovative Graduate Study Program for Materials Research, Information, and Technology (MERIT-WINGS) of the University of Tokyo and JSPS KAKENHI Grant No. JP24KJ0857.
@@
Supplementary Notes for: Error Crafting in Probabilistic Quantum Gate Synthesis
S1 Formalism of probabilistic synthesis under diamond distance
@@
In the main text, we have discussed the probabilistic synthesis that minimizes the diamond distance $d_{\diamond}(\Phi_{1},\Phi_{2})$ between two unitary channels $\Phi_{1}$ and $\Phi_{2}$. Here, we provide the formal definition of the diamond distance, which is related with the diamond norm as $d_{\diamond}(\Phi_{1},\Phi_{2}):=\frac{1}{2}\|\Phi_{1}-\Phi_{2}\|_{\diamond}$ where $\Phi_{1},\Phi_{2}$ are linear maps from $\mathcal{L}(\mathcal{H}_{1})$ to $\mathcal{L}(\mathcal{H}_{2})$ ($(\mathcal{L}(\mathcal{H}_{i}))$ : the entire set of linear map on $\mathcal{H}_{i}$) for an $d$-dimensional Hilbert space $\mathcal{H}(=\mathcal{H}_{1}=\mathcal{H}_{2})$:
@@
Here, $\|X\|_{1}={\rm Tr}[\sqrt{X^{\dagger}X}]$ denotes the trace norm, ${\rm Id}_{d}$ is an identity channel acting on $d$-dimensional Hilbert space, and $\mathbb{D}(\mathcal{H}_{1}\otimes\mathcal{H}_{2})$ is the set of physical density matrices on $\mathcal{H}_{1}\otimes\mathcal{H}_{2}$. It is known that the diamond norm $\|\Phi\|_{\diamond}$ can be efficiently calculated via semidefinite programming (SDP) for general linear maps as [58, 59]
@@
Now let us consider probabilistic synthesis of a target CPTP map $\hat{\Upsilon}$ using a set of CPTP implementable maps $\{\Upsilon_{j}\}$ under the following minimization problem:
@@
S1.1 Simplification of probabilistic synthesis in single-qubit case
Although SDP is efficient in a sense that there exists a polynomial-time algorithm with respect to the problem matrix size, in practice the runtime grows quite rapidly, and hence it is desirable to simplify the minimization problem. For the single-qubit case, the calculation of diamond distance boils down to that of trace distance between Choi matrices. While one can show $\frac{1}{d}\|\Phi_{1}-\Phi_{2}\|_{\diamond}\leq\frac{1}{d}\|\mathcal{J}(\Phi_{1})-\mathcal{J}(\Phi_{2})\|_{1}\leq\|\Phi_{1}-\Phi_{2}\|_{\diamond}$ for general $d$-dimensional completely positive maps $\Phi_{1}$ and $\Phi_{2}$, for the single-qubit mixed unitary channel case it is known to saturate the lower bound as $\|\Phi_{1}-\Phi_{2}\|_{\diamond}=\|\mathcal{J}(\Phi_{1})-\mathcal{J}(\Phi_{2})\|_{1}/2.$ Therefore, the remnant error of probabilistic synthesis, $\min_{p}\frac{1}{2}\left\|\hat{\Upsilon}-\sum_{j}p_{j}\Upsilon_{j}\right\|_{\diamond}$, can now be computed by the following simpler semidefinite programming problem:

### v2

@@ Results
If one chooses to employ the PEC method as the QEM method for FTQC, it suffices to guarantee only the feasibility under the Pauli constraint. Indeed, the PEC method is powerful in the sense that one may perform unbiased estimation when the error channel is precisely known. However, we find that the remnant errors can be counteracted much more effectively by utilizing the fact that they can be characterized completely by classical computation. For instance, we may wish to craft local depolarizing errors, since many limits and performance of quantum algorithms, as well as QEM methods, are analyzed under the assumption of the local depolarizing noise Aharonov et al. (1996); Koczor (2021a, b); Quek et al. (2022); Cai et al. (2023). We may alternatively wish to craft the remnant errors to be some biased Pauli noise. We find that this is extremely beneficial for the case of the Pauli rotation gates. Namely, we can perform error detection or correction in the logical layer, which reduces the sampling overhead to the extent that QEM methods designed for general channels cannot achieve.
@@
First, let us discuss the crafting of the depolarizing channel. Following a similar strategy as in Theorem 1, we can mathematically prove that there exists a set of shift unitaries $\{\mathcal{V}^{c\epsilon}_{j}\}$ such that we can achieve quadratic suppression of error under the depolarizing constraint, with the sole difference being that the number of shift unitaries is 9 instead of 7. The formal statement and the proof are provided in Supplementary Note S7. Our numerical simulation also shows similar favorable properties as seen in the case of Pauli constraint (See Supplementary Note S3).
@@
Next, let us consider crafting the remnant error to be biased. In particular, we focus on the Pauli Z rotation gate and aim to bias the error to be XY error, since it can be detected using a single ancilla qubit with less sampling overhead than the lower bound for QEM methods that do not rely on mid-circuit measurement Tsubouchi et al. (2023) (concrete circuit structure provided in Fig. 3). While we have not found a guaranteed set of shift unitaries, we find that the nine shift unitaries for the depolarizing noise are of practical use, if we allow violation of the constraint as $p_{z}\leq 0.01p$ where $p\coloneqq p_{x}+p_{y}+p_{z}$ (see Supplementary Notes S3 for details).
@@
The success of error crafting is highlighted in Fig. 4. As can be seen from Fig. 4(a), while the rate of Pauli errors $(\frac{p_{x}}{p},\frac{p_{y}}{p},\frac{p_{z}}{p})$ is distributed quite homogeneously if we do not impose any constraint on the error profile, we find that the constraints can be reflected with high accuracy. We verify that $p_{x}=p_{y}=p_{z}$ is satisfied up to the order of machine precision for the depolarizing constraint, while for the XY-biased noise, we find that we cannot completely eliminate the Z error; there is a small residual component of $p_{z}=O(\epsilon^{3})$ even if we allow the XY-biased noise to consist of non-unital components. We also attempted to craft X-biased noise, while it did not succeed in most instances.
@@
Figure 3: Circuit structure for (a) error detection and (b) logical error correction for Pauli Z rotation gate $R_{z}$ with remnant error of $\mathcal{E}_{\rm rem}$. We may further combine with the PEC method, for which the Pauli operation $\mathcal{P}$ with dotted square is executed virtually via updating the Pauli frame. Note that the feedback operation in (b) is also done via the Pauli frame.
Figure 4: Numerical analysis on error crafting for remnant errors. Here we display the distribution of error rate ratio $(q_{x},q_{y},q_{z})\coloneqq(\frac{p_{x}}{p},\frac{p_{y}}{p},\frac{p_{z}}{p})$ on a plane $q_{x}+q_{y}+q_{z}=1$, where $p_{a}\coloneqq\chi_{aa}~{}(a\in\{x,y,z\})$ is the rate of each Pauli channel in $\mathcal{E}_{\rm rem}$ and $p\coloneqq p_{x}+p_{y}+p_{z}=O(\epsilon^{2})$ corresponds to the total error rate. Note that here we show results whose violation of the constraints are below threshold values (see Supplementary Notes S3 for details). The unitary synthesis is done for $\epsilon=10^{-4}$ with shift factor $c=2.0,5.0$ for depolarizing and XY constraints, respectively.
Error crafting using CPTP maps
While we have considered the error crafting using synthesized unitaries $\{U_{j}\}$, we can also apply the framework to CPTP maps. As a demonstration, here we consider the target unitary $\hat{U}$ to be a Pauli Z rotation, and constitute the set of implementable CPTP maps $\{\Lambda_{j}\}$ as the channels of synthesized unitaries accompanied with entangled measurement and feedback as shown in Fig. 3(b). We have numerically found that, the remnant error can be suppressed cubically small as $\epsilon^{3}$, as opposed to the quadratic suppression realized in the error crafting for unitaries. Since the optimal unitary synthesis yield T-count of $3\log_{2}(1/\epsilon)+O(1)$, such an improvement allows us to synthesize a Pauli rotation gate with a T-count of $\log_{2}(1/\varepsilon)+O(1)$ on average. Compare this with the scaling of $1.5\log_{2}(1/\varepsilon)$ realized by mixed synthesis, which achieve quadratic improvement in $\varepsilon$ (see Fig. 5). We guide the readers to the Supplementary Note S4S5 for a detailed discussion on the search for appropriate $\Lambda_{j}$.
@@
Gate Synthesis Crafted? Remnant error QEM method Sampling overhead Comment
@@
SU(2) Mixed Unitary $\epsilon^{2}$ not Pauli PEC $1+4\epsilon^{2}$ Pauli frame unavailable
SU(2) Mixed Unitary $\epsilon^{2}$ Pauli/depol PEC $1+4\epsilon^{2}$
SU(2) Mixed Unitary $\epsilon^{2}$ Pauli/depol Rescaling $1+2\epsilon^{2}$
$R_{z}$ Unitary $\epsilon$ coherent PEC $1+4\epsilon$ Pauli frame unavailable
@@
$R_{z}$ Mixed Unitary $\epsilon^{2}$ not Pauli PEC $1+4\epsilon^{2}$ Pauli frame unavailable
$R_{z}$ Mixed Unitary $\epsilon^{2}$ Pauli/depol/XY PEC $1+4\epsilon^{2}$
$R_{z}$ Mixed Unitary $\epsilon^{2}$ Pauli/depol/XY ED + PEC $1+(q_{x}+q_{y}+4q_{z})\epsilon^{2}$ $q_{z}\sim\epsilon$ in numerics
$R_{z}$ Mixed Unitary $\epsilon^{2}$ Pauli/depol/XY ED + Resc. $1+(q_{x}+q_{y}+2q_{z})\epsilon^{2}$ $q_{z}\sim\epsilon$ in numerics
$R_{z}$ Mixed CPTP $\epsilon^{3}$ Pauli Z PEC $1+4\epsilon^{3}$ Single ancilla
$R_{z}$ Mixed CPTP $\epsilon^{3}$ Pauli Z Rescaling $1+2\epsilon^{3}$ Single ancilla
@@
Table 2: Sampling overhead of error countermeasures under synthesis error of $\epsilon$. The remnant error indicates the diamond distance $d(\mathcal{E}_{\rm rem},\mathcal{I})$ where $\mathcal{I}$ is the identity channel. We assume that unitary gate synthesis is done with an accuracy of $\epsilon$, and compare the performance when we further perform mixed or error-crafted synthesis. The sampling overhead $\gamma$ is indicated for a single gate, and the total overhead is multiplicative as $\prod_{l}\gamma_{l}$ when overhead is $\gamma_{l}$ for the $l$-th gate. The rescaling technique assumes that the underlying quantum circuit is sufficiently deep so that the white noise (WN) approximation holds Dalzell et al. (2024); Morvan et al. (2024); Tsubouchi et al. (2024). We assume that the white noise approximation holds when the rescaling technique is applied. The cubic suppression in the last two rows is verified numerically up to $\epsilon^{3}=10^{-9}$ as shown in Fig. 5. Note that applying PEC for crafted and uncrafted synthesis differs significantly. We cannot rely on the Pauli frame in the latter, and an actual single-qubit operation is required.
@@
Figure 5: Trade-off between T-count and synthesis accuracy for Pauli rotation gates. The red, yellow, and blue regions indicate the achievable accuracy without any sampling overhead when we employ unitary, mixed, and error-crafted synthesis. The blue circles indicate the error solely from crafted synthesis for $R_{z}(\theta)$ with $\theta\in\{\frac{m\pi}{32}\}_{m\neq 0~{}{\rm mod4}}$ under various synthesis accuracy. Note that total error combined with hardware noise is bounded by the chained black line, which indicates the contribution from noise in T gates. Here we choose error rate of $10^{-11}$ per T gate. The dotted red, yellow, and green lines are given by the empirical T-count of $3\log_{2}(1/\varepsilon)$, $1.5\log_{2}(1/\varepsilon)$, and $\log_{2}(1/\varepsilon)$, respectively, and the real lines indicate the total error.
@@
Even if the Z error cannot be eliminated completely, such an error crafting drastically alters the sampling overhead. By performing the error detection operation as shown in Fig. 3(a), we can mitigate Pauli X and Y errors with sampling overhead of $\gamma_{\rm XY}=1+p_{x}+p_{y},$ which is even below the lower bound of QEM sampling overhead that does not rely on mid-circuit measurement Tsubouchi et al. (2023). While the Z error is undetectable and hence we shall rely on other QEM methods such as PEC or rescaling with overhead of $\gamma_{Z}=1+\alpha p_{z}$, the total sampling overhead is $\gamma_{\rm tot}=\gamma_{\rm XY}\gamma_{Z}\sim 1+p_{x}+p_{y}+\alpha p_{z}\sim 1+\epsilon^{2}$. When we perform QEM for $L$ gates, the total sampling overhead $\gamma_{\rm tot}^{L}$ is quartically small compared to that of PEC.
An intuitive understanding of the reduction of the T-count can be obtained from a geometric interpretation. Let us first briefly describe the standard volume argument Harrow et al. (2002) for SU(2) synthesis to yield scaling of $3\log_{2}(1/\varepsilon)+O(1)$, and then proceed to explain the improved scaling of $\log_{2}(1/\varepsilon)+O(1)$. We can interpret the problem of designing a synthesis algorithm, i.e., ensuring any $d$-dimensional unitary to be approximated by accuracy of $\varepsilon$, as packing $\epsilon$-ball into the manifold of SU($d$). The volume of $\varepsilon$-ball in $(d^{2}-1)$-dimensional space is $O(\varepsilon^{d^{2}-1}),$ in particular $V=O(\varepsilon^{3})$ for SU(2). Hence, in order to cover the entire space, we need at least $O(1/V)=O(1/\varepsilon^{3})$ unique unitaries. Note that any Clifford+T gate can be expressed uniquely by the canonical form proposed by Matsumoto and Amano Matsumoto and Amano (2008). Specifically, the number of unique Clifford+T operator with T-count of $t$ is $192\cdot(3\cdot 2^{t}-2)$. We therefore deduce that the lower bound of T-count scales as $3\log_{2}(1/\varepsilon)+O(1)$. It has been numerically verified that this bound is indeed nearly saturated by a direct search algorithm Morisaki et al. (2024), which reflects the fact that the grid points corresponding to Clifford+T unitary are distributed homogeneously within the manifold so that the gate set quickly mimics the Haar unitary ensemble Haferkamp et al. (2023).
When we perform the mixed CPTP synthesis, since we can correct coherent error of X rotation via the feedback operation, we may allow it to be arbitrarily large. This effectively projects out one of the dimensions of the manifold to be covered. Correspondingly, the volume of the ball is regarded as $V=O(\varepsilon^{2})$. From the volume argument, this leads to the T-count of $2\log_{2}(1/\varepsilon)+O(1)$ for the synthesis of unitaries to be included in the mixture. By combining with the fact that appropriate mixed synthesis allows quadratic suppression of error and hence reduces the T-count by a factor of two, we arrive at the scaling of $\log_{2}(1/\varepsilon)+O(1)$.
@@
Let us briefly comment on the classical precomputation time of crafted synthesis. Our results in Fig. 5 exploits the direct search algorithm for SU(2) gate synthesis whose computational complexity scales as $O(1/\epsilon)$ on average Morisaki et al. (2024). Considering that currently known unitary synthesis algorithms for $R_{z}$ gates based on number-theoretic approach Kliuchnikov et al. (2015); Ross and Selinger (2016) or repeat-until-success circuits Bocharov et al. (2015); Kliuchnikov et al. (2023) yield nearly optimal gate sequence with computational complexity of $O({\rm polylog}(1/\epsilon))$, the complexity of direct search algorithms seems to be a significant overhead. Meanwhile, we observe that the actual computation time per synthesis is observed to be milliseconds for the accuracy regime we have investigated. Considering that such computations can be embarrassingly parallelized, and also that number of rotation gates for early FTQC circuits is of $10^{4}$ for quantum dynamics simulation Campbell (2021); Beverland et al. (2022) and $10^{6}$ for ground state energy simulation Kivlichan et al. (2020), we envision that classical computation time is not bottleneck for such applications.
@@
Last but not least, while we have exclusively discussed the error and sampling overhead only for the synthesis error, it must be kept in mind that, in reality, there is also incoherent noise that indirectly limits the synthesis accuracy. As can be seen from Fig. 5, one shall aim for the sweet spot where the total error is minimized. While we have displayed the case when we have designed the FTQC to contribute the majority of logical error to magic state distillation, we may flexibly change the scheme to introduce errors in Clifford operations as well.
Discussion
@@
Numerous future directions are envisioned. First, it is essential to seek whether error crafting is possible for synthesis of multiqubit unitaries. Combined with the fact that the error can potentially be suppressed exponentially smaller with the qubit count for multiqubit case Akibue et al. (2024), it is crucial to extend the framework proposed in this work to perform practical quantum computation. Second, it is intriguing in terms of resource theoretic argument to assess the probabilistic transformability between channels. While this has been studied for state transformation in the context of magic state or entanglement distillation Regula (2022), the argument on quantum channels is left as an interesting open question. Third, it is an intriguing open question whether there is a lower bound on the computational complexity to perform error-crafted synthesis.
Data availability
Data and codes for the work are available via GitHub repository Yoshioka (2024).
@@ Acknowledgement
We thank fruitful discussions with Suguru Endo, Ali Javadi-Abhari, Yuuki Tokunaga, and Kunal Sharma. This work was supported by JST CREST Grant Number JPMJCR23I4, JST Moonshot R&D Grant Number JPMJMS2061, MEXT Q-LEAP Grant No. JPMXS0120319794, and No. JPMXS0118068682. N.Y. wishes to thank JST PRESTO No. JPMJPR2119, JST ERATO Grant Number JPMJER2302, JST Grant Number JPMJPF2221, and the support from IBM Quantum. S.A. is supported by JST, PRESTO Grant no.JPMJPR2111. This work is supported by MEXT Quantum Leap Flagship Program (MEXT Q-LEAP) Grant No. JP- MXS0118067394 and JPMXS0120319794, JST COI- NEXT Grant No. JPMJPF2014. K.T. is supported by World-leading Innovative Graduate Study Program for Materials Research, Information, and Technology (MERIT-WINGS) of the University of Tokyo and JSPS KAKENHI Grant No. JP24KJ0857.
Author contributions
N.Y. conceived the idea. N.Y., S.A., and K.T. performed theoretical analysis. N.Y. and H.M. developed the numerical codes. All authors contributed to scientific discussion and writing of the paper.
Competing Interests
Y.S. and S.A. own stock/options in Nippon Telegraph and Telephone. Y.S. owns stock/options in QunaSys Inc.
@@
Supplementary Notes for: Error Crafting in Mixed Quantum Gate Synthesis
S1 Limitation in applying twirling for non-Clifford operation
In the main text, we have explained using Table 1 that neither Pauli or Clifford twirling is not available for non-Clifford gates in general. Here we provide details on this argument.
Let $\mathcal{U}$ denote a non-Clifford single-qubit gate and $\mathcal{N}$ denote a noise channel associated to it. When we wish to realize the implementation of the target unitary with twirled noise, its channel representation can be written as
$\mathbb{E}_{\mathcal{D}}[\mathcal{D}^{\dagger}\circ\mathcal{N}\circ\mathcal{D}]\circ\mathcal{U},$ (S1)
where $\mathcal{D}$ is averaged over some single-qubit unitary subgroup such as the Pauli or Clifford group. However, a gate operation is always accompanied by the noise as $\mathcal{N}\circ\mathcal{U}$, and we cannot insert $\mathcal{D}$ in between.
One of the options is to modify the expression as
$\mathbb{E}_{D}[\mathcal{D}^{\dagger}\circ\mathcal{N}\circ\mathcal{U}\circ(\mathcal{U}^{\dagger}\circ\mathcal{D}\circ\mathcal{U})],$ (S2)
which requires us to perform the following in sequence:
  1. (1)
    Operate conjugated unitary $UDU^{\dagger}$ without noise.
  2. (2)
    Operate noisy target unitary.
  3. (3)
    Operate $D$ without noise.
While the second and third steps are commonly considered to be achievable (coined as “hard” and “easy” operations in literature Wallman and Emerson (2016)), the problem is in the first step, since $UDU^{\dagger}$ is a non-Clifford operation in general. For instance, if we consider $U=e^{i\theta Z}$ and $D=X$, we must implement $UDU^{\dagger}=e^{i\theta Z}Xe^{-i\theta Z}=e^{2i\theta Z}X.$ Since we assume that rotation gates are noisy, this cannot be implemented without injecting additional untwirled noise. This is however completely contrary to what we desire to do.
An exception is when the target gate belongs to the third level of the Clifford hierarchy. In this case, $UDU^{\dagger}$ is a Clifford operation if $D$ is a Pauli operator, and hence we can perform Pauli twirling by using Clifford gates. However, the Clifford twirling is not available; if we take the Hadamard gate, we can show that
$(THT^{\dagger})\cdot Z\cdot(THT^{\dagger})^{\dagger}=THZHT^{\dagger}=TXT^{\dagger}=SX.$ (S3)
This indicates that $THT^{\dagger}$ does not map a Pauli operator into a Pauli operator, and hence not a Clifford gate.
One might consider another style of twirling by implementing $D^{\dagger}UD$ instead of $UDU^{\dagger}$. This direction is based on the fact that Eq. (S1) can also be formally rewritten as
$\mathbb{E}_{\mathcal{D}}[\mathcal{D}^{\dagger}\circ\mathcal{N}\circ(\mathcal{D}\circ\mathcal{U}\circ\mathcal{D}^{\dagger})\circ\mathcal{D}].$ (S4)
The requirement for this formulation is (1) the noise channel of $D^{\dagger}UD$ is all identically $\mathcal{N}$ irrespective of $\mathcal{D}$ Wallman and Emerson (2016), and (2) the noise in $\mathcal{D}$ is negligible. When the target is $U=e^{i\theta P}$ with $P\in\{X,Y,Z\}$ and $\theta$ being the rotation angle, the requirement is that noise channel for Pauli rotation gates with angle $\theta$ are identical to each other. However, this is unlikely in general, since the noise channel in unitary synthesis is unitary; to ensure the noise channel to be identical, we must accurately perform the unitary synthesis to control the synthesis error, which is beyond the actual task of synthesizing a target unitary up to accuracy of $\epsilon$.
S2 Formalism of mixed synthesis under diamond distance
@@
In the main text, we have discussed the mixed synthesis that minimizes the diamond distance $d_{\diamond}(\Phi_{1},\Phi_{2})$ between two unitary channels $\Phi_{1}$ and $\Phi_{2}$. Here, we provide the formal definition of the diamond distance, which is related with the diamond norm as $d_{\diamond}(\Phi_{1},\Phi_{2}):=\frac{1}{2}\|\Phi_{1}-\Phi_{2}\|_{\diamond}$ where $\Phi_{1},\Phi_{2}$ are linear maps from $\mathcal{L}(\mathcal{H}_{1})$ to $\mathcal{L}(\mathcal{H}_{2})$ ($(\mathcal{L}(\mathcal{H}_{i}))$ : the entire set of linear map on $\mathcal{H}_{i}$) for an $d$-dimensional Hilbert space $\mathcal{H}(=\mathcal{H}_{1}=\mathcal{H}_{2})$:
@@
Here, $\|X\|_{1}={\rm Tr}[\sqrt{X^{\dagger}X}]$ denotes the trace norm, ${\rm Id}_{d}$ is an identity channel acting on $d$-dimensional Hilbert space, and $\mathbb{D}(\mathcal{H}_{1}\otimes\mathcal{H}_{2})$ is the set of physical density matrices on $\mathcal{H}_{1}\otimes\mathcal{H}_{2}$. It is known that the diamond norm $\|\Phi\|_{\diamond}$ can be efficiently calculated via semidefinite programming (SDP) for general linear maps as Watrous (2009, 2012)
@@
Now let us consider mixed synthesis of a target CPTP map $\hat{\Upsilon}$ using a set of CPTP implementable maps $\{\Upsilon_{j}\}$ under the following minimization problem:
@@
S2.1 Simplification of mixed synthesis in single-qubit case
Although SDP is efficient in a sense that there exists a polynomial-time algorithm with respect to the problem matrix size, in practice the runtime grows quite rapidly, and hence it is desirable to simplify the minimization problem. For the single-qubit case, the calculation of diamond distance boils down to that of trace distance between Choi matrices. While one can show $\frac{1}{d}\|\Phi_{1}-\Phi_{2}\|_{\diamond}\leq\frac{1}{d}\|\mathcal{J}(\Phi_{1})-\mathcal{J}(\Phi_{2})\|_{1}\leq\|\Phi_{1}-\Phi_{2}\|_{\diamond}$ for general $d$-dimensional completely positive maps $\Phi_{1}$ and $\Phi_{2}$, for the single-qubit mixed unitary channel case it is known to saturate the lower bound as $\|\Phi_{1}-\Phi_{2}\|_{\diamond}=\|\mathcal{J}(\Phi_{1})-\mathcal{J}(\Phi_{2})\|_{1}/2.$ Therefore, the remnant error of mixed synthesis, $\min_{p}\frac{1}{2}\left\|\hat{\Upsilon}-\sum_{j}p_{j}\Upsilon_{j}\right\|_{\diamond}$, can now be computed by the following simpler semidefinite programming problem: